Mauro Products distributes a single product, a woven basket whose selling price is $27 per unit and whose variable expense is $19 per unit. The company’s monthly fixed expense is $10,400.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

Selling price per unit = $27

Variable cost per unit = $19

Fixed cost = $10,400

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 27 - 19

= $8

1.

Break even point in units = Fixed cost/Contribution margin per unit

= 10,400/8

= 1,300

2.

Break even point sales dollar = Break even point in units x Selling price per unit

= 1,300 x 27

= $35,100

3.

Fixed expenses increased by $600

New fixed expense = 10,400 + 600

= $11,000

New Break even point in units = New fixed expense/Contribution margin per unit

= 11,000/8

= 1,375

New Break even point sales dollar = New Break even point in units x Selling price per unit

= 1,375 x 27

= $37,125

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.